Calculate the gears for cutting metric threads of following

pitch. i) 4mm pitch, ii) 5.25 mm pitch. The lead screw of lathe

contains 6 TPI. The lathe is supplied with 20 to 120 teeth in step

of 5and an additional gear wheel of having 127 teeth.


1
Expert's answer
2021-06-04T16:08:02-0400

For metric threads,

"\frac{Driver \space teeth}{Driver \space teeth}=\frac{5mp}{127}"

where, n= number of threads per inch i.e TPI

P= Pitch of the thread to be cut

(i) 4mm

"\frac{Driver \space teeth}{Driven \space teeth}=\frac{5mp}{127} =\frac{5*6*4}{127}=\frac{120}{127}Ans"

The gear train will consist of 120 teeth on the spindle gear and 127 teeth on the lead screw

"5.25mm" Pitch


"\frac{Driver \space teeth }{Driven \space teeth}=\frac{5mp}{127} =\frac{5*6*5.25}{127}=\frac{5*6\frac{21}{4}}{127}"

"=\frac{105*6}{4*127}=\frac{105}{40}*\frac{60}{127} A ns"

Numerators and denominators are multiplying by 10

For 5.25mm pitch, compound gear train to be used with 105 teeth on spindle gear and 127 teeth on the intermediate gear , 60 teeth intermediate gear drives a 40 teeth gear on the lead screw.